40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 632     Entry time: Thu Jul 3 16:18:51 2008
Author: rob 
Type: Summary 
Category: Locking 
Subject: specgrams 
I used ligoDV to make some spectrograms of DARM_ERR (1), QPDX (2), and QPDY (3). These show the massive instability from 30-40Hz growing in the XARM in the last two minutes of a reasonably high power lock (arm powers up to 30). It's strange that it only shows up in one arm.

CARM is on PO-DC, for both the MCL and the AO path.
DARM is on AS166Q.
Attachment 1: darm_specg.png  67 kB  | Hide | Hide all
darm_specg.png
Attachment 2: qpdx_specg.png  59 kB  | Hide | Hide all
qpdx_specg.png
Attachment 3: qpdy_specg.png  61 kB  | Hide | Hide all
qpdy_specg.png
ELOG V3.1.3-